This page shows the status of our Belnet Network and Services. Currently these are updated manually.
To report incidents, please contact our Belnet Service Desk at 02 790 33 00 or servicedesk@belnet.be.

All systems are operational

Events

14th January 2020

Backbone Network issue impacting some of our customers

After a change implementation on some optical switches, we are currently experiencing a technical issue on our network which affect connectivity for some of our customers. Our teams are currently investigating to restore connectivity and our services . Belnet customers can contact our Service Desk at 02 790 33 00. We apologize for the inconvenience.

Workaround is now also in place for our secondary router in the PoP of Kortrijk. This means that all affected services on the network are restored at this moment. We keep working towards a general solution with the vendor at a high momentum.

At 03:05, a workaround has been put in place to restore IP connectivity to the main router in our PoP in Kortrijk. We are in the process of implementing the workaround for the secondary router in Kortrijk as well.

We are performing some emergency changes on the optical network, based on the troubleshooting with the vendor. This will impact optical services very briefly but is mandatory for the proceedings of our troubleshooting. The change has been tested and will create an outage of approx. 3 minutes.

In the meantime several optical services have been reestablished. Belnet teams are still working to restore remaining unavailable services in POP Kortrijk.

Recovery process is still on going for POP Kortrijk and optical services. Focus is to restore these services as soon as possible.

A workaround has been installed in POP Brugge and Hasdie and IP services are restored in these locations. Optical services and POP Kortrijk recovery are still ongoing.

The configuration recovery test is still ongoing. We will update you as soon as possible.

We are testing the configuration recovery of optical switches in a test environment. If it goes fine we will implement it in production and will have news around 14:00

Optical switch cards with problems have been restarted. Belnet teams are now checking and restoring the configuration on these optical switch cards. Investigations on other possible resolution paths are continuing.

All Belnet teams continue investigating the problem will all our ressources on all possible resolution path in collaboration with the third party. Engineers are on site. We do not have an estimate to resolution yet. We will keep you updated within the next hour.

Incident is escalated to our Optical Supplier higher level of Support. Remote troubleshooting hasn't helped, a cold restart of the cards did not work. We have sent engineers on site to take additional actions on our optical switches.

Impacted customers are customers connected in our POP's of Kortrijk, Brugge and Hasselt

11th January 2020

Louvain-La-Neuve Power outage LLN

Due to a power outage in LLN, some of our customers connected to LLN have lost their connectivity to the Belnet network. Our teams are going on site to manage the equipment when the power will be restored. Belnet customers can contact our Service Desk at 02 790 33 00. We apologize for the inconvenience.

23rd December 2019